Solve for exact answer lnx = 11
Answer by Theo(3504)

ln(x) = 11

you want to find x?

In general,

ln(x) = y if and only if e^y = x

if y = 11, this means that:

e^11 = x

e^11 = 59874.14171 which means that:

x = 59874.14171

This means that:

ln(x) = 11 becomes:

ln(59874.14171) = 11 which becomes:

11 = 11 confirming the value of x is good.